Use the Laplace transform to solve the given initial-value problem.

y' + 3y = e5t,  y(0) = 2

y(t) =


1
Expert's answer
2021-04-14T14:48:46-0400

y' + 3y = e5t,  y(0) = 2


Taking Laplace transform on both sides


"L(y')+3L(y) = L(e^{5t})"

"sY(s)-y(0)+3Y(s) = \dfrac{1}{(s-5)}"

"sY(s)-2+3Y(s) = \dfrac{1}{(s-5)}"

"Y(s)(s+3) - 2 = \dfrac{1}{(s-5)}"


"Y(s) = \dfrac{2}{(s+3)} + \dfrac{1}{(s+3)(s-5)}"

"Y(s) = \dfrac{(2s-9)}{(s+3)(s-5)}"

Resolving into partial fraction we get,


"\dfrac{(2s-9)}{(s+3)(s-5)} = \dfrac{A}{(s+3)} + \dfrac{B}{(s-5)}"

"(2s-9) = A(s-5) + B(s+3)"

We get, "A = \dfrac{15}{8}"

"B = \dfrac{1}{8}"

Hence, we get


"Y(s) = \dfrac{15}{8(s+3)} + \dfrac{1}{8(s-5)}"

Now, taking the inverse Laplace Transform

We get,

"y(t) = \dfrac{15(e^{-3t})}{8} + \dfrac{e^{5t}}{8}"